(x-9)(x-9)=90
(x+9)(x+9)=90

Answer by mananth(16946)

(x-9)(x-9)=90
+ / - (x-9)= sqrt90
+(x-9)=9.48
x1= 18.48
-(x-9)= 9.48
-x+9= = 9.48
-x= 9.48-9
-x2= 0.48
x2= - 0.48
x1 & x2 are the roots of the equation